OpenStudy (anonymous):

(9!)(4!)/(7!)(2!) Evaluate

OpenStudy (anonymous):

3456

OpenStudy (anonymous):

u kno google can insta tell u these answers

OpenStudy (anonymous):

\[\frac{ 9! \times 4! }{ 7! \times 2! }\] \[\frac{ 9 \times 8 \times 7 \times 6 \times 5 \times 4 \times 3 \times 2 \times 1 \times 4 \times 3 \times 2 \times 1 }{ 2 \times 1 \times 7\times 6\times 5\times 4\times 3\times 2\times 1 }\]

OpenStudy (anonymous):

cancel out numbers.

OpenStudy (anonymous):

and you get 864

OpenStudy (anonymous):

@hahd ur wrong my man. @SpecialMathStudent i would double check your work if i were you.

OpenStudy (anonymous):

you probably used https://www.wolframalpha.com/input/?i=%289!%29%284!%29+%2F+%287!%29%282!%29

OpenStudy (anonymous):

lol im not wrong the way he arranged the brackets im perfectly correct

OpenStudy (anonymous):

i got 864 also

OpenStudy (anonymous):

took me a lil bit but i got it..

OpenStudy (anonymous):

yeah me too

OpenStudy (anonymous):

thanks guys for the help

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

@hahd there are no brackets? Only parenthesis.

OpenStudy (anonymous):

and its not confusing?

OpenStudy (anonymous):

lol ok

OpenStudy (unklerhaukus):

\[\frac{\color{teal}{9!}\times\color{orange}{4!}}{\color{brown}{7!}\times\color{cornflowerblue}{2!}}=\frac{\color{teal}{9\times8\times7!}\times\color{orange}{4\times3\times2!}}{\color{brown}{7!}\times\color{cornflowerblue}{2!}}=\color{teal}{9\times8}\times\color{orange}{4\times3}\]

OpenStudy (anonymous):

@UnkleRhaukus lol that is what i did :P

OpenStudy (unklerhaukus):

Yes, same method, but it's a little simpler looking if you dont expand all of the factorials , because you know they are gonna cancel anyways,

OpenStudy (anonymous):

yeah but i wanted to show the user who answered the question like step by step, just in case he wanted to see where i got all the numbers. :)

OpenStudy (anonymous):

who asked*

OpenStudy (anonymous):

nice colorful text though ;) lol

OpenStudy (unklerhaukus):

you method is better for someone who dosent understand the notation very well, so i am very glad you used it, i was just showing the simplest way to solve, once you do understand. i probably wouldn't have left my answer as is, if your full expansion wasn't there already. [the solutions are complementary]

OpenStudy (anonymous):

haha yeah i usually don't even expand that long, i either do what you do, or just use my calc. :D

OpenStudy (anonymous):

But good job @UnkleRhaukus :)